Tripling the greater of two consecutive even integers gives the same result as subtracting 10 from the lesser even integer. What are the integers? I need help on everything in this questing I need to know how to do it and show my work! Please help
Found 2 solutions by wuwei96815, glabow:
Answer by wuwei96815(245)

The first problem is translating the word problem into a math problem.

Let n = one of the even integers Let n+2 = the other integer

Now let us write an equation and solve it:

3(n+2) = n - 10 This equation expresses what is written in the word problem.

3(n+2) = n - 10
3n + 6 = n - 10
3n - n = -10 -6
2n = -16
n = -8
n+2 = -6

You can check these answers by substituting them into the original equation.

Answer by glabow(165)

First, what are two consecutive even integers. If you let x be the first one, what would the next even integer be? If you said x+2 you would be right.
So, the smaller of the two is x, the larger is x+2.
Now you know that 3 times the larger one is equal to the smaller one less 10.
3 times the larger one is written algebraicly like this: 3(x+2).
The smaller one less 10 is written like this: x - 10.
You know 3(x+2) = x-10.
Can you solve this equation for x?
3x+6 = x-10 [distributive rule]
2x+6=-10 [subtract x from both sides]
2x=-16 [subtract 6 from both sides]
x=-8 [divide both sides by 2]
x+2=-6
So three times the greater number (-6) is equal to the smaller number (-8) minus 10.
[Checking: 3(-6)=-18. -8-10=-18.]